OpenStudy (anonymous):

Solve the fractions for z: z/z-3 + z/z^2-9 = z+4/z+3

jimthompson5910 (jim_thompson5910):

\[\frac{z}{z-3}+\frac{z}{z^2-9}=\frac{z+4}{z+3}\] \[\frac{z}{z-3}+\frac{z}{(z-3)(z+3)}=\frac{z+4}{z+3}\] \[z(z+3)+z=(z+4)(z-3)\] \[z^2+3z+z=z^2+z-12\] \[z^2+3z+z-z^2-z+12=0\] \[3z+12=0\] \[3z=-12\] \[z=\frac{-12}{3}\] \[z=-4\] I'll let you perform the check.

OpenStudy (anonymous):

z = -4 is correct